Project

General

Profile

09/08/2014

Notes taken by M.Tamsett & C.Group. Docdb:
http://nova-docdb.fnal.gov:8080/cgi-bin/DisplayMeeting?conferenceid=1782

Attendance

Nate, Michael, Jeny, Chris, Matthew, Pavan, Craig, Dominick, Kirk, Satish, Kanika, Jonathan, Gavin

News (Matthew)

Simulation aren’t quite ready yet, they’re still waiting on some ND stuff. We expect we will hear from them this week.

Dataset definitions from make_sim_fcl (Nate)

Nate proposes that we define core dataset definitions in make_sim_fcl and then use these to build datasets for downstream tiers using the “isparentof” or “ischildof” functionality. Everyone agreed that this was a sensible idea. Nate will use the small datasets being produced for Simulation validation as a test case. He also pointed out that make_sim_fcl isn’t executable from the “first-ana” branch. Naming conventions are coded into this and will follow the standard patterns:

prod_DATATIER_RELEASE_DETECTOR_TRIGGERSTREAM_SPECIAL
prod_DATATIER_RELEASE_DETECTOR_FLAVORSET_SPECIAL

Naming conventions and Utilities/batch cleanup discussion (all)

Margaret, Andrew, others are interested in how we version control and validate our production scripts. Consensus was that we should move to a new package for “production” tools which only contains in use scripts.

Action Item: Matthew will set up this package, including underscores in names for clarity.

Calibration keep-up handover (Satish, Paola)

Keep-up is running smoothly again now. Satish and Paola will meet to organise handover this week. It was mentioned that we need to add ND into this and that we should probably include the tricell and/or activity triggers.

Action Item: Matthew will contact calibration to work out which streams they need.

Reco keep-up update (Jeny, Dominick)

This is now ready to go and is working in development. They’re just waiting on the new art tag. The plan is to start with ND data, then follow with the FD.

Baseline samples for geo-jittered/crygenie comparisons (Nate & Gavin)

Both the baseline and offset samples are through Reco and PID and we expect CAFs to be ready in the next day or so.

Art update (Kanika)

Everything working with the exception of some issues/inconsistencies with checksum match for classes that changed, but these aren’t vital and we can turn them off and move forward if need be. She wants to run some test jobs before we can sign off and tag and then will be ready on the time scale of a couple of day. Kanika intends to make two tags – new art, and first ana.

Mr-XX (Kanika)

The reco jobs are finished and she will soon submit jobs for pid. It was noted that LEM still needs stuff changing to work - module labels. Matthew has contacted Duyang to find out the status/plans of MR-brem.

Postgres tests (Dominick)

Dominick ran some tests last week and will continue to run some more this week.

CPN locks (Chris)

Chris reports that the rate defining step in his jobs can be CPN locks. He enquired about the use of the pnfs scratch area.

Action Item: Craig will follow up on this with regards to how we can use it.

AOB

  • Gavin raised the question of user facilities to put files into sam? and where do the lines lie between us and users need to plan this with the collaboration? This needs follow up with a larger audience.
  • Jeny reported that she’s seeing raw2root delays in getting files to tape and that she wants to do audit of past data.
  • R.Hatcher request for detailed statistics particularly for the genie step. Nate will work with him and Eric to get these together.